The nos of silicon atoms / m^3is 5 * 10 ^28  ,  this is doped simultaneously with 5 * 10^22 atoms/m^3 of Arsenic and 5*10^20 per m^3 of Indium

Calculate the nos of holes and electron

Given n i  =  1.5 * 10^16 / m^3

Ans ne = 4.95 * 10^22 , nh = 4.75 * 10^9

Please give detailed solution i am not getting it


Grade:10

1 Answers

Pushkar Aditya
10 years ago
Under thermal equilibrium the product of free negative and positive charge concentrations is constant quantity given by nenh=ni^2. Arsenic is an n type impurity and Indium is a p type impurity. No. of electrons=5*10^22-5*10^20=4.95*1022 m^-3. So we get nh= 4.75 * 10^9 using the formula nenh=ni^2.